OpenStudy (anonymous):

A social networking site recorded that 80% of the people who became members of the site one month were between 30 and 40 years old. The equation below represents this information, where x represents the total number of people who became members of the site that month. x = 0.8x + 72 How many people who became members of the site that month were between 30 and 40 years of age

OpenStudy (fibonaccichick666):

so you want to solve for x. any ideas on your first step?

OpenStudy (anonymous):

no

OpenStudy (fibonaccichick666):

ok so let's try this\[x=.8x+72\] now can you get the x terms on the same side?

OpenStudy (fibonaccichick666):

sorry try this \[3x=4x+5\]

OpenStudy (anonymous):

dont you do this: subtract 5 on both sides and then after that divide?

OpenStudy (fibonaccichick666):

not quite, you want to get only x terms on the left(or right) and only non x terms on the other side. can you tell me which terms are 'x-terms'?

OpenStudy (anonymous):

3x and4x

OpenStudy (fibonaccichick666):

good so if you want to isolate x (aka get all the x terms together on one side) what do you need to do?

OpenStudy (anonymous):

subtract 3x from 4x

OpenStudy (fibonaccichick666):

can you show me this please?

OpenStudy (anonymous):

its going to be x which is basically 1x

OpenStudy (fibonaccichick666):

but can you write the equation/ show me your work; otherwise, I cannot see if you did it correctly

OpenStudy (fibonaccichick666):

ex. \[6x=4x\]\[6x-4x=0\]\[2x=0\]

OpenStudy (anonymous):

3x=4x+5 -3x -3x 1x=5 x=5

OpenStudy (anonymous):

right?

OpenStudy (fibonaccichick666):

well not quite if you subtract 3x from both sides as you did you will get\[0=x+5\]

OpenStudy (fibonaccichick666):

do you see why?

OpenStudy (anonymous):

no

OpenStudy (goformit100):

First of all for solving this question do you know the conception of : Theory of equations.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok good night this is to much for my brain

OpenStudy (fibonaccichick666):

uh ok